40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14620     Entry time: Fri May 17 17:01:08 2019     In reply to: 14615
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY suspension characterization 

To investigate my mapping of the eigenfrequencies to eigenmodes, I checked the Oplev spectra for the last few hours, when the Oplev spot has been on the QPD (but the optic is undamped).

  1. Based on Attachment #1, I can't figure out which peak corresponds to what motion.
    • The most prominent peak (judged by peak height) is at 0.771 Hz for both PITCH and YAW
    • Assuming the peak at 0.92 Hz is the other angular mode, the PIT/YAW decoupling is poor in both peaks, only ~factor of 2 in both cases.
  2. Why are the POS and SIDE resonances sensed so asymmetrically in the PIT and YAW channels? There's a factor of 10 difference there...

So, while I conclude that my first-contact residue removal removed a constraint from the system (hence the pendulum dynamics are accurate and there are 6 eigenmodes), more thought is needed in judging what is the appropriate course of action.

Attachment 1: etmy_oplevs.pdf  11 kB  Uploaded Fri May 17 18:01:59 2019  | Hide | Hide all
etmy_oplevs.pdf
ELOG V3.1.3-